### v1

Abstract
The Gaussian process (GP) is a widely used probabilistic machine learning method for stochastic function approximation, stochastic modeling, and analyzing real-world measurements of nonlinear processes. Unlike many other machine learning methods, GPs include an implicit characterization of uncertainty, making them extremely useful across many areas of science, technology, and engineering. Traditional implementations of GPs involve stationary kernels (also termed covariance functions) that limit their flexibility and exact methods for inference that prevent application to data sets with more than about ten thousand points. Modern approaches to address stationarity assumptions generally fail to accommodate large data sets, while all attempts to address scalability focus on approximating the Gaussian likelihood, which can involve subjectivity and lead to inaccuracies. In this work, we explicitly derive an alternative kernel that can discover and encode both sparsity and nonstationarity. We embed the kernel within a fully Bayesian GP model and leverage high-performance computing resources to enable the analysis of massive data sets. We demonstrate the favorable performance of our novel kernel relative to existing exact and approximate GP methods across a variety of synthetic data examples. Furthermore, we conduct space-time prediction based on more than one million measurements of daily maximum temperature and verify that our results outperform state-of-the-art methods in the Earth sciences. More broadly, having access to exact GPs that use ultra-scalable, sparsity-discovering, nonstationary kernels allows GP methods to truly compete with a wide variety of machine learning methods.
keywords:
Machine learning, Gaussian processes, Positive definite functions, Nonstationarity, Gridded data products
1 Introduction
Gaussian processes (GPs) are a preeminent framework for stochastic function approximation, statistical modeling of real-world measurements, and non-parametric and nonlinear regression within machine learning (ML) and surrogate modeling. One reason for the popularity of GPs is that they are analytically tractable and can be fully specified in terms of a prior mean function and a kernel, also known as a covariance function. Additionally, unlike most ML methods, GPs are “probabilistic” in the sense that they naturally include measures of uncertainty in predictions and, when implemented in a Bayesian setting, can furthermore account for and propagate uncertainty in their inferred hyperparameters. GPs have been broadly used across many areas of science, technology, and engineering, including forward modeling [1, 2, 3], uncertainty quantification [4, 5], and autonomous experimentation [6, 7, 8, 9].
However, there are two main reasons why GPs are not used more broadly for analyzing modern data sets. First, traditional implementations of GPs involve stationary kernels [e.g., the squared exponential kernel is widely used in Gaussian process applications; 10] that are too inflexible to model real-world data [11]. The fact that standard stationary kernels are so widely used leads to the general impression that GPs themselves are inflexible, as opposed to a particular implementation of GPs being inflexible. Second, evaluating the Gaussian likelihood for $N$ data points scales with $\mathcal{O}(N^{2})$ in memory and $\mathcal{O}(N^{3})$ in computation, such that GPs are computationally infeasible for as few as several thousand data points. Given the fact that modern data sets quite commonly involve millions or even billions of data points, in an ML context, this lack of scalability leads users to gravitate towards neural networks and deep learning methods that are easily applied to big data.
The statistics and ML literature contains a wide variety of methods to address these two primary limitations of GPs. Much work has been done to derive and implement nonstationary kernels, wherein the dependence between two input locations cannot be written as a function of their separation distance [see, e.g., 12, 13, 14, 15, 16, 17]. However, with a few exceptions, these methods are generally difficult to implement for large data sets due to their large numbers of hyperparameters and the need for specialized algorithms for training. These methods can also lead to overfitting; multiple authors [18, 19, 20] have pointed out the potential dangers of using over-parameterized models (such as GPs with nonstationary kernels) without regularization techniques such as posterior averaging [21] or prior regularization [22, 23] in Bayesian non-parametric models [24, 3, 25].
Regarding scalability, a variety of methods have been proposed that derive approximations to the GP likelihood, e.g., local GP experts [26], inducing-point and low-rank methods [27, 28, 29], structure-exploiting methods [30], and Vecchia approximations [31, 32, 33]. We refer the interested reader to [34] for a direct comparison of many of these methods on a common data set. These so-called “approximate GP” methods are not, in general, agnostic to the underlying kernel, and theoretical properties are, in most cases, derived only for stationary kernels. Furthermore, each of these methods requires multiple subjective choices – e.g., Vecchia approximations necessitate specifying the number of neighbors to condition on, the ordering of points, and the selection of nearest neighbors – that may introduce sensitive ad hoc choices in a given analysis.
A final limitation of most kernels used in the literature (both stationary and nonstationary) is that they cannot model covariances that are exactly zero. In almost all cases, a kernel will decay to zero as the separation between two points increases; however, the kernel will still return covariances that are technically nonzero, even for points that are very far apart in the input space. It is our hypothesis that most (if not all) real-world data sets implicitly include at least some degree of sparsity in the covariance, but typical classes of kernels cannot flexibly model both zero and nonzero covariances. Compactly-supported kernels [see, e.g., 35, 36, 37, 38, 39] can model both zero and nonzero covariances in a data-driven manner, albeit usually with stationary kernels. The approximate GP methods also “model” sparsity, but the implied sparsity is fully determined by ad hoc choices such as the ordering of points and neighbor selection. The combination of nonstationary and also compactly supported kernels would enable more appropriate modeling of real-world data and yield computational shortcuts for the matrix operations needed to calculate the Gaussian likelihood. In short, the density of the GP prior covariance is not fundamental to the method itself but a result of poorly chosen kernels. Instead of using these standard kernels and then retroactively working on scalability, we propose to give the GP the ability to discover sparsity during the model fitting process. A demonstration of our proposed methodology is shown in Figure 1.
Figure 1: Ground truth function over $[0,10]$ with 50 noisy measurements (panel a.); the synthetic function is both sparse and nonstationary. Panel (b) shows the resulting posterior mean function with 95% uncertainty intervals for four methods: a stationary Gaussian process with a Matérn (non-sparse) kernel, the Vecchia approximation (a leading approximate GP method), and our sparsity-discovering kernel combined with both stationary and nonstationary core kernels. Panel (c) shows posterior median covariance estimates (gray areas indicate where the kernel is identically zero). Our kernel has three clear benefits: first, it can discover sparsity in the data while using an exact Gaussian likelihood; second, it can infer the nonstationary structure in the data wherein diagonal elements of the covariance are larger at the endpoints of the domain; third, it outperforms exact and leading approximate methods in terms of root mean square error (RMSE) and continuous rank probability score (CRPS; lower scores for both metrics indicate a better fit).
In this paper, we introduce a novel kernel that explicitly demonstrates how to apply the proof-of-concept methodology proposed in [40] to real-world data sets. The overall kernel consists of a compactly supported kernel that can discover and encode sparsity combined with an arbitrary “core” kernel, yielding ultra-flexible kernels that allow one to apply exact nonstationary GPs to arbitrarily large data sets. Together with prior regularization, the induced sparsity also prevents overfitting and leads to a better interpretation of the data since it preserves the flexibility of a nonstationary kernel. We also propose a complete Bayesian stochastic model based on the new kernel that accounts for a generic prior mean function and measurement error and allows posterior prediction for unobserved input locations with uncertainty quantification. While high-performance computing (HPC) resources are needed to apply our method to truly massive data sets, we also show how the proposed kernel can be used to conduct fully Bayesian analysis for moderately sized data sets with limited computational resources, e.g., using a personal laptop. The efficacy of our novel kernel is demonstrated using both synthetic data examples and a real-world data set from the Earth sciences.
2 Sparsity-discovering nonstationary kernels
One of the core ideas underpinning this methodology is that real-world data sets almost always involve sparse covariances, but typical classes of kernels cannot flexibly model zero covariances. Let $y({\bf x})$ be a mean-zero Gaussian process defined on a generic input space $\mathcal{X}\subset\mathbb{R}^{d}$ with a kernel (also called a covariance function) $C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})\equiv Cov\big{(}y({\bf x}),y({\bf x}^{\prime})\big{)}$ for all ${\bf x},{\bf x}^{\prime}\in\mathcal{X}$ that describes the covariance between the process $y({\bf x})$ and is known up to a set of hyperparameters ${\boldsymbol{\theta}}_{y}$. In order to explicitly model sparsity in the stochastic process $y(\cdot)$, we propose modeling the kernel as
$C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})=C_{\text{core}}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{\text{core}})\times C_{\text{sparse}}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{\text{sparse}}),$
where $C_{\text{core}}$ is a positive definite kernel (stationary or nonstationary), $C_{\text{sparse}}$ is a sparsity-inducing kernel (defined below in Eq. 2), and ${\boldsymbol{\theta}}_{y}=({\boldsymbol{\theta}}_{\text{core}},{\boldsymbol{\theta}}_{\text{sparse}})$. So long as $C_{\text{core}}$ and $C_{\text{sparse}}$ are positive definite, their product will be positive definite [41], yielding a powerful framework for building flexible kernels.
2.1 Nonstationary kernels that discover sparsity
One way to specify $C_{\text{sparse}}$ such that it is positive definite and can model sparsity is via a kernel that utilizes sums and products of so-called “bump functions”
$g({\bf x})=\left\{\begin{array}[]{ll}a\exp\left\{b\left[1-(1-||\mathbf{x}-\mathbf{h}||^{2}/r^{2})^{-1}\right]\right\}&\text{if }||\mathbf{x}-\mathbf{h}||^{2}<r\\ 0&\text{else},\end{array}\right.$ (1)
where $a\geq 0$ represents the amplitude, $b>0$ is a shape parameter (see Figure G.1 for a visualization of the effect of $b$), $\mathbf{h}\in\mathbb{R}^{d}$ is the centroid, and $r>0$ is the radius [42]. Note that $g({\bf x})$ is smooth and exactly zero for point-to-centroid distances greater than or equal to $r$, which is what enables the covariance matrix to be sparse. We combine these bump functions to yield a sparse kernel
$C_{\text{sparse}}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{\text{sparse}})=s_{0}f_{0}({\bf x},{\bf x}^{\prime};r_{0})+\sum_{i=1}^{n_{1}}f_{i}({\bf x})f_{i}({\bf x}^{\prime}).$ (2)
In Eq. 2, each $f_{i}~{}\forall i\geq 1,$ is the sum of bump functions $g_{ij}$ from Eq. 1 (each with hyperparameters $\{{\bf h}_{ij},a_{ij},b_{ij},r_{ij}\}$), i.e.,
$f_{i}({\bf x})=\sum_{j=1}^{n_{2}}g_{ij}({\bf x}),\hskip 17.22217pti=1,\dots,n_{1},$
$f_{0}({\bf x},{\bf x}^{\prime};r_{0})$ is a compactly supported stationary kernel, and $s_{0}$ is a scalar parameter to adjust the relative magnitude of $f_{0}({\bf x},{\bf x};r_{0})$ and the amplitudes $\{a_{ij}\}$. We include an additive stationary kernel in Eq. 2 because there might otherwise be cases where $C({\bf x},{\bf x};{\boldsymbol{\theta}}_{\text{sparse}})=0$; specifying the stationary kernel to be compactly supported ensures that the resulting covariance matrix is still sparse. Among the broad range of possibilities for specifying a compactly supported $f_{0}$ [see, e.g., 36, 37, 38, 39], we opt to use members of the Wendland family of polynomials [35], e.g.,
$f_{0}({\bf x},{\bf x}^{\prime};r_{0})=\left\{\begin{array}[]{ll}\left(1-\frac{||{\bf x}-{\bf x}^{\prime}||}{r_{0}}\right)^{8}\left(35\left(\frac{||{\bf x}-{\bf x}^{\prime}||}{r_{0}}\right)^{3}+25\left(\frac{||{\bf x}-{\bf x}^{\prime}||}{r_{0}}\right)^{2}+\frac{8||{\bf x}-{\bf x}^{\prime}||}{r_{0}}+1\right)&\text{if }||{\bf x}-{\bf x}^{\prime}||^{2}<r_{0},\\ 0&\text{else.}\end{array}\right.$
The choice of $f_{0}$ has implications on the differentiability of the resulting kernel; we return to this point later in the section. Considering $n_{1}$ and $n_{2}$ fixed and denoting the hyperparameters of the $j^{\text{th}}$ bump function $g_{ij}$ of the $i^{\text{th}}$ component of Eq. 2 as $\{{\bf h}_{ij},a_{ij},b_{ij},r_{ij}\}$, the entire vector of hyperparameters for $C_{\text{sparse}}$ is ${\boldsymbol{\theta}}_{\text{sparse}}=\left(s_{0},r_{0},\{{\bf h}_{ij},a_{ij},b_{ij},r_{ij}:i=1,\dots,n_{1};j=1,\dots,n_{2}\}\right)$ which involves a total of $2+n_{1}n_{2}(d+3)$ hyperparameters (recall $d$ is the dimension of the input space, such that ${\bf h}\in\mathcal{X}\subset\mathbb{R}^{d}$). While we suppose that $n_{1}$ and $n_{2}$ are fixed a priori, these are of course important quantities for a given implementation of the kernel; in a Bayesian framework, we use suitable prior distributions to regularize the sparse kernel, effectively enabling the user to choose arbitrary values of $n_{1}$ and $n_{2}$ such that the data can then determine the optimal number of bump functions (see Section 3.1 for further details).
In order to show how the various components of $C_{\text{sparse}}$ in Eq. 2 combine to flexibly model sparsity in the overall kernel $C_{y}$, we present an illustrative example. For this example, we set $n_{1}=2$ and $n_{2}=3$ and take $\mathcal{X}=[0,1]$ with $N=100$ equally spaced locations (hence $d=1$). Figure 2(a) shows the bump functions $g_{ij}(x)$ and their sums $f_{i}(x)$ using fixed values of the bump function hyperparameters. Starting with the leftmost column of Figure 2, the product of the sums-of-bump-functions $\sum_{i}f_{i}(x)f_{i}(x^{\prime})$ in the middle row reveals the effect of the bump functions: for a given $i$, the support (i.e., domain of nonzero values) of the bump functions define distance-unrelated parts of the domain that are correlated; furthermore, across the various $i$, a given point in the input space can have a nonzero correlation with different sets of points that does not depending on the distances from this point to these sets. For example, note that both because $f_{1}(0.17)>0$ and $f_{1}(0.7)>0$, $C_{\text{sparse}}(0.17,0.7)>0$ even though these points are relatively far apart. Furthermore, because $f_{2}(0.17)>0$ and $f_{2}(0.9)>0$ (but $f_{2}(0.7)=0$), we have $C_{\text{sparse}}(0.17,0.9)>0$ but $C_{\text{sparse}}(0.7,0.9)=0$. While the radii $r_{ij}$ specify regions of zero versus nonzero correlation, the amplitudes $a_{ij}$ and shapes $b_{ij}$ determine the relative magnitude of off-diagonal nonzero correlations and how sharply peaked the correlations are near the bump function centers, respectively shown in Figure 2(b) and (c). Finally, increasing the radii in Figure 2(d) predictably decreases the overall sparsity of $C_{\text{sparse}}$ and (together with relatively small $b$) increases the areas of high correlations. The compactly supported stationary covariance $f_{0}(x,x^{\prime};r_{0}=0.15)$ serves to ensure that the diagonal elements of $C_{\text{sparse}}$ are nonzero, with $r_{0}$ determining the extent to which points very close together are correlated within the sparse kernel.
(a) Baseline example         (b) Shape×$\times 0.1$             (c) Amplitude×$\times 2$             (d) Radius×$\times 1.5$
Bump functions + sums
$\sum_{i=1}^{2}f_{i}(x)f_{i}(x^{\prime})$
$C_{\text{sparse}}(x,x^{\prime})$
Figure 2: Illustration of how the bump functions are used to define the sparse covariance $C_{\text{sparse}}$ when $n_{1}=2$ and $n_{2}=3$ for $N=100$ equally spaced locations on the unit interval $[0,1]$. The top row shows the bump functions $g_{ij}$ and their sum $f_{i}$ for each $i=1,2$ on the unit interval; the middle row shows image plots of the resulting matrices of entries calculated from the product of bump function sums $\sum_{i}f_{i}(x)f_{i}(x^{\prime})$; the bottom row shows image plots of $C_{\text{sparse}}$, i.e., the normalized sum $f_{0}(x,x^{\prime};0.15)+\sum_{i}f_{i}(x)f_{i}(x^{\prime})$. The red (solid) and blue (dashed) lines in the middle row correspond to the effect of bump functions for $i=1$ and $i=2$, respectively. The different columns show the effects of groups of hyperparameters for fixed bump function centroids: starting from the leftmost column, we consecutively change the shape parameters $b$ (column b.), the amplitudes $a$ (column c.), and the radii (column d.).
In summary, each subset of hyperparameters in $C_{\text{sparse}}$ serves an important purpose for modeling zero and nonzero correlations: the centroids $\{{\bf h}_{ij}\}$ determine portions of the domain that have distance-unrelated (nonstationary) nonzero correlations; the radii $\{r_{ij}\}$ determine the size of the domain subsets that have distance-unrelated nonzero correlations; the shapes $\{b_{ij}\}$ determine how sharply peaked the distance-unrelated nonzero correlations are (sharply peaked for large $b$ and flat for small $b$); the amplitudes $\{a_{ij}\}$ determine the relative magnitude of distance-unrelated nonzero correlations (larger than average $a$ yields larger than average off-diagonal nonzero correlations and vice versa); the compactly supported kernel length-scale $r_{0}$ determines scales at which nearby points are correlated; and the scaling factor $s_{0}$ determines the relative magnitude of the compactly supported kernel and the bump functions.
Our specification of $C_{\text{sparse}}$ implies a nonstationary kernel, irrespective of the choice of $C_{\text{core}}$: note that we cannot write Eq. 2 as a function of the separating vector ${\bf x}-{\bf x}^{\prime}$ since the bump functions in Eq. 1 cannot be written as a function of ${\bf x}-{\bf x}^{\prime}$. Our approach explicitly models sparsity directly in the covariance matrix to capture the inherent sparsity of the dependence structure in real-world datasets; alternatively, approximate methods [e.g., inducing points; 43, 29] focus on approximating the GP with a subset of the data to achieve computational savings. In other words, both classes of methods leverage sparsity to achieve computational savings, but the sparsity implied by our method is data driven and describes the underlying structure in the data set of interest.
Since the bump functions are infinitely differentiable [42], we now consider the choice of $f_{0}$ on the differentiability of $C_{y}$ at the origin.
Proposition 1.
The kernel $C_{y}$ has $w=\min\{w_{0},w_{\text{core}}\}$ continuous derivatives at zero, where $f_{0}$ has $w_{0}$ continuous derivatives at the origin and $C_{\text{core}}$ has $w_{\text{core}}$ continuous derivatives at zero.
The proof is provided in Appendix A.
2.2 Modeling choices for the core kernel
In principle, there are no constraints regarding statistical modeling of the core kernel; any positive semi-definite function can be used. In contrast to many approximate methods to scale GPs, our framework allows us to use the full suite of tools that have been developed for kernel modeling, including stationary and nonstationary kernels. For example, even using standard stationary kernels (e.g., the Matérn class) as the core kernel results in the overall kernel being both nonstationary and sparse (since $C_{\text{sparse}}$ is itself nonstationary). Later in the paper, we utilize parametric families of nonstationary kernels [drawn from, e.g., 15, 44, 45]; in principle, however, one could use any other nonstationary kernel including deep kernels [17] or deep GPs [16]. In the absence of specific knowledge regarding the form of the core kernel, “parametric” nonstationary kernels [46] such as those derived in [15] and later extended in [45] have been shown to perform well in practice. See Appendix B for more details.
2.3 Special cases
Before concluding this section, we briefly describe three special cases of our kernel that can be learned from data via training. First, $C_{y}$ is sparse and nonstationary by construction, whether or not the core kernel is nonstationary. However, three other cases are possible:
  1. 1.
    $C_{y}$ becomes sparse and stationary if $C_{\text{core}}=1$ and all bump functions are set equal to zero (possible if either the radii $\rightarrow 0$ or all amplitudes are zero).
  2. 2.
    $C_{y}$ becomes non-sparse and stationary if $r_{0}\rightarrow\infty$ and $C_{\text{core}}$ is either (1) chosen to be stationary or (2) collapses to stationary through regularization.
  3. 3.
    $C_{y}$ becomes non-sparse and nonstationary if $s_{0},r_{0}\rightarrow\infty$ and $C_{\text{core}}$ is nonstationary.
Prior specification and regularization allow the kernel to learn the underlying true state (see Appendix D.2).
3 Bayesian stochastic modeling and inference
We now propose a Bayesian framework for statistical modeling with a Gaussian process (Section 3.1) along with an approach to posterior prediction with uncertainty quantification (Section 3.2). Details regarding the computational requirements of our method are discussed in Section 3.3, and a pair of synthetic data tests conducted to verify the performance of our methodology (described in Appendix E) are summarized in Section 3.4.
3.1 Bayesian GP with prior specification, regularization, and training
Let $\{z({\bf x}):{\bf x}\in\mathcal{X}\}$ be the observed value of a univariate stochastic process on a given Euclidean domain $\mathcal{X}\subset\mathbb{R}^{d}$. A general framework for modeling $z({\bf x})$ with a GP is
$z({\bf x})=y({\bf x})+\varepsilon({\bf x}),$
where $\varepsilon(\cdot)$ is a stochastic error component that is independently distributed as ${N}(0,\tau^{2}({\bf x}))$, with $\varepsilon(\cdot)$ and $y(\cdot)$ are independent. We furthermore assume that the error variance process $\tau^{2}(\cdot)$ is known up to hyperparameters ${\boldsymbol{\theta}}_{z}$. The stochastic process $y(\cdot)$ is modeled as a Gaussian process, where the prior mean function $E[y({\bf x})]={\bf w}({\bf x})^{\top}{\boldsymbol{\beta}}$ is linear in a set of covariates or basis functions and the kernel $C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})\equiv Cov\big{(}y({\bf x}),y({\bf x}^{\prime})\big{)}$ is assumed known up to hyperparameters ${\boldsymbol{\theta}}_{y}$ and describes the covariance between the process $y(\cdot)$ for all ${\bf x},{\bf x}^{\prime}\in\mathcal{X}$.
For a fixed and finite set of $N$ input locations $\{{\bf x}_{1},...,{\bf x}_{N}\}\in\mathcal{X}$, the Gaussian process assumptions imply that the random (observed) vector ${\bf z}=\left[z({\bf x}_{1}),...,z({\bf x}_{N})\right]^{\top}$ has a multivariate Gaussian distribution $p({\bf z}|{\bf y},{\boldsymbol{\theta}}_{z})={N}\big{(}{\bf y},{\bf\Delta}({\boldsymbol{\theta}}_{z})\big{)}$, where ${\bf\Delta}({\boldsymbol{\theta}}_{z})=diag[\tau^{2}({\bf x}_{1}),\dots,\tau^{2}({\bf x}_{N})]$ and ${\bf y}=\left[y({\bf x}_{1}),...,y({\bf x}_{N})\right]^{\top}$. Conditional on the other hyperparameters in the model, the vector ${\bf y}$ is distributed as $p({\bf y}|{\boldsymbol{\beta}},{\boldsymbol{\theta}}_{y})={N}\big{(}{\bf W}{\boldsymbol{\beta}},{\bf\Omega}({\boldsymbol{\theta}}_{y})\big{)}$, where ${\bf W}=[{\bf w}({\bf x}_{1})^{\top},\dots,{\bf w}({\bf x}_{N})^{\top}]^{\top}$ and the elements of ${\bf\Omega}({\boldsymbol{\theta}}_{y})$ are $\Omega_{ij}\equiv C_{y}({\bf x}_{i},{\bf x}_{j};{\boldsymbol{\theta}}_{y})$. To simplify training, one can integrate over the process $y(\cdot)$ to arrive at the marginal distribution of the data:
[…]
where ${\boldsymbol{\theta}}=({\boldsymbol{\theta}}_{z},{\boldsymbol{\theta}}_{y})$. The kernel for the marginalized process is
$C_{z}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}})=C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})+\tau({\bf x})\tau({\bf x}^{\prime})I_{\{{\bf x}={\bf x}^{\prime}\}},\hskip 17.22217pt\text{for all }{\bf x},{\bf x}^{\prime}\in\mathcal{X},$ (4)
where $I_{\{\cdot\}}$ is an indicator function. Our approach to modeling $C_{y}$ was described in Section 2.
In a Bayesian setting, one must define prior distributions for the unknown mean and covariance parameters $p({\boldsymbol{\beta}},{\boldsymbol{\theta}})$; all inference for ${\boldsymbol{\beta}}$ and ${\boldsymbol{\theta}}$ is then based on the marginalized posterior for these parameters conditional on ${\bf z}$:
$p({\boldsymbol{\beta}},{\boldsymbol{\theta}}|{\bf z})\propto p({\bf z}|{\boldsymbol{\beta}},{\boldsymbol{\theta}})p({\boldsymbol{\beta}},{\boldsymbol{\theta}}).$ (5)
Without specific prior knowledge, it is convenient to assume a priori that the prior distribution $p({\boldsymbol{\beta}},{\boldsymbol{\theta}})=p({\boldsymbol{\beta}})p({\boldsymbol{\theta}})$ (i.e., ${\boldsymbol{\beta}}$ and ${\boldsymbol{\theta}}$ are statistically independent) and furthermore that $p({\boldsymbol{\beta}})=N_{p}({\bf 0},k{\bf I}_{p})$, where $k$ is a large constant (e.g., $k=100^{2}$) such that the prior for the mean coefficients is noninformative. Regarding prior distributions for ${\boldsymbol{\theta}}_{z}$ and ${\boldsymbol{\theta}}_{y}=({\boldsymbol{\theta}}_{\text{core}},{\boldsymbol{\theta}}_{\text{sparse}})$, proper but noninformative priors will be specified for the elements of ${\boldsymbol{\theta}}_{z}$. Specific choices for ${\boldsymbol{\theta}}_{\text{core}}$ depend of course upon the specific core kernel; see Appendix B for more details.
For ${\boldsymbol{\theta}}_{\text{sparse}}$, without loss of generality we fix the bump function shapes $\{b_{ij}\}$ to be 1; otherwise, we specify proper but noninformative priors for the compact stationary length scale and bump function radii:
$s_{0}\sim U(0,10^{5});\hskip 21.52771ptr_{0}\sim U(0,D_{0});\hskip 21.52771ptr_{ij}\sim U(0,D_{r}),\hskip 8.61108pti=1,\dots,n_{1};j=1,\dots,n_{2}$
where $U(a,b)$ denotes a uniform distribution on the interval $(a,b)$. Here, $D_{0}$ and $D_{r}$ are user-specified upper limits on the local influence of the Wendland and the size of the bump functions’ influence, respectively. The prior for the bump function locations $\{{\bf h}_{ij}\}$ is then $h_{ij}^{k}\sim U(l_{k},u_{k})$, where the input space $\mathcal{X}\subset\mathbb{R}^{d}$ (assuming a Euclidean domain) has coordinate-specific lower and upper bounds $\{l_{k},u_{k}:k=1,\dots,d\}$ (here, $h_{ij}^{k}$ is the $k^{\text{th}}$ element of ${\bf h}_{ij}$). Similar priors can be defined for non-Euclidean input spaces. Finally, we propose a simple regularization prior for the bump function amplitudes $\{a_{ij}\}$, where we assume probabilistically that $a_{ij}\sim\text{Bernoulli}(\pi_{ij})$. In other words, the amplitudes are binary random variables that allow individual bump functions $g_{ij}$ to be turned on ($a_{ij}=1$) or off ($a_{ij}=0$), where $\pi_{ij}=\text{Pr}(a_{ij}=1)$. Lastly, the probabilities $\pi_{ij}$ are assigned uniform priors on the unit interval. In the case that the $ij^{\text{th}}$ bump function is unnecessary, this formulation will prevent it from having unintended influence on the data-driven sparsity.
For GPs with trained hyperparameters, the posterior distribution (5) is not available in closed form regardless of prior choice, and so we must resort to Markov chain Monte Carlo (MCMC) methods to conduct inference on ${\boldsymbol{\beta}}$ and ${\boldsymbol{\theta}}$ [47]. See Appendix C for further details.
3.2 Posterior prediction with uncertainty quantification
Posterior prediction of the process $y(\cdot)$ for either the observed inputs or a distinct set of $N_{P}$ unobserved inputs $\{{\bf x}^{*}_{1},...,{\bf x}^{*}_{N_{P}}\}\in\mathcal{X}$ is straightforward given the Gaussian process assumptions used here. In a Bayesian setting, one can conduct either conditional simulation (i.e., conditioning on the data) or unconditional simulation (i.e., without conditioning on the data). Define ${\bf y}_{P}=({\bf y},y({\bf x}^{*}_{1}),...,y({\bf x}^{*}_{N_{P}}))$; for conditional simulation the predictive distribution of interest in the Bayesian setting is
[…]
Based on the Gaussian assumptions, $p({\bf y}_{P}|{\boldsymbol{\beta}},{\boldsymbol{\theta}},{\bf z})$ can be calculated in closed form as $N\left({\bf m}_{{\bf y}|{\bf z}},{\bf C}_{{\bf y}|{\bf z}}\right)$, where
[…]
A Monte Carlo estimate of $p({\bf y}_{P}|{\bf z})$ can then be obtained using draws from the posterior, i.e., $p({\bf y}_{P}|{\bf z})\approx\sum_{l}p({\bf y}_{P}|{\boldsymbol{\beta}}^{(l)},{\boldsymbol{\theta}}^{(l)},{\bf z})$, where $\{{\boldsymbol{\beta}}^{(l)},{\boldsymbol{\theta}}^{(l)}\}$ are draws from the posterior $p({\boldsymbol{\beta}},{\boldsymbol{\theta}}|{\bf z})$. For unconditional simulation, one can obtain samples of ${\bf y}$ using posterior draws via ${\bf y}^{*}={\bf W}_{P}{{\boldsymbol{\beta}}^{(l)}}+{\bf U}^{(l)}{\bf o}$, where ${\bf U}^{(l)}$ is a Cholesky decomposition of ${\bf C}_{\bf y}$ (which is a function of posterior draw ${{\boldsymbol{\theta}}}_{l}$) and the elements of ${\bf o}$ are iid $N(0,1)$. In other words, this yields ${\bf y}^{*}\sim N({\bf W}_{P}{{\boldsymbol{\beta}}^{(l)}},{\bf C}^{(l)}_{\bf y})$.
3.3 Computational considerations
In order to utilize $C_{z}$ within a Bayesian framework, each iteration of the Markov chain Monte Carlo for the posterior Eq. 5 requires at least one evaluation of the log of the probability density function given in Eq. 3 (the log-likelihood), i.e., $\log p({\bf z}|{\boldsymbol{\beta}},{\boldsymbol{\theta}})\propto-0.5\log\det({\bf C}_{z})-0.5({\bf z}-{\bf W}{\boldsymbol{\beta}})^{\top}{\bf C}_{z}^{-1}({\bf z}-{\bf W}{\boldsymbol{\beta}})$, where the elements of ${\bf C}_{z}$ come from Eq. 4. From a computational standpoint, there are three main tasks associated with this evaluation: (1) calculating the covariance matrix ${\bf C}_{z}$, (2) calculating the log-determinant $\log\det({\bf C}_{z})$, and (3) solving the quadratic form $({\bf z}-{\bf W}{\boldsymbol{\beta}})^{\top}{\bf C}_{z}^{-1}({\bf z}-{\bf W}{\boldsymbol{\beta}})$. The sparsity-discovering nature of our kernel means there are no shortcuts for calculating the covariance matrix for a fixed set of hyperparameters: different elements of ${\bf C}_{z}$ will be zero vs. nonzero for different hyperparameter configurations. As such, we must first calculate ${\bf C}_{z}$ in dense form in batches that are large enough to utilize the chosen hardware well. Computationally, the covariance batches are dense but contain many zeros because of our kernel designs; therefore, each batch is cast to a sparse format on the worker before being transferred back to the host process. There, the full covariance matrix is assembled and exists only in a sparse format leading to minimal RAM requirements. For particularly large datasets, the computation of the batches can be performed by utilizing distributed high-performance-computing architecture. For example, [40] applied a related kernel to a climate data set of more than 5 million points using HPC resources at the National Energy Research Scientific Computing Center at Lawrence Berkeley National Laboratory.
Steps two and three, namely the log-determinant calculation and the linear-system solve related to the quadratic form can be performed at low computational costs due to the sparse covariance matrix. For example, using the Lanczos method to estimate the log-determinant has a time complexity of $\mathcal{O}(km+k^{2})$, where $m$ is the number of nonzero entries in the matrix and $k$ is the number of iterations, which depends on the eigenvalue distribution of the matrix. Further, the conjugate gradients methods to solve linear systems has time complexity $\mathcal{O}(mk)$, where $m$ is the number of nonzero elements in the matrix and $k$ is the number of iterations which depend on the condition number of the matrix. For poorly conditioned covariance matrices, we use the minimum residual method which has the same time complexity but is less sensitive to the condition number. In practice, we frequently see sparsity numbers of $1/10000$ for large problems of 1 million data points which, together with the proposed kernel design, leads to log-determinant and linear solve compute times in the order of a few second. For smaller problems ($<100,000$), we can take advantage of LU or Cholesky decompositions.
The fact that our methodology yields no shortcuts for calculating ${\bf C}_{z}$ is in sharp contrast with approximate methods, e.g., the nearest neighbor Gaussian process (NNGP) proposed in [48] and the Vecchia approximations proposed in [32]. First, the NNGP and Vecchia approximations imply a sparse precision matrix (i.e., inverse covariance), while our method implies a sparse covariance. Second, the implicit conditioning structure of each method implies that for a given ordering of points and method for choosing nearest neighbors [see choices C1-C5 given in Section 2.3 of 32], some elements of the precision matrix will always be nonzero, and other elements will always be exactly zero. The fixed nonzero elements are known ahead of time (again, based on specific choices made for ordering of points and selecting neighbors), such that likelihood evaluations for NNGP and Vecchia never involve dense matrices. At the end of the day, our methodology is still an exact Gaussian process: some of the usual challenges associated with computation for a GP cannot be avoided. However, what we gain is substantially less subjectivity: no a priori choices need to be made regarding, e.g., the number of neighbors to condition on, the ordering of points, and the selection of nearest neighbors.
3.4 Synthetic data tests
Before applying our methodology to a real-world data set, we conducted two synthetic data tests to ensure that our proposed kernel is robust to mis-specification of the number of bump functions and performs well relative to competing methods across a variety of ground-truth data states.
Our first test (see Appendix D for complete details) verifies that our proposed sparse kernel can correctly identify the bump function placement and size, and furthermore, the regularization method for the bump functions described in Section 3.1 effectively eliminates the effect of any bump functions that may be unnecessary. For a real data example, this leads us to recommend specifying more bump functions than one might think is truly necessary (i.e., err on the side of making $n_{1}$ and $n_{2}$ large). Of course, more bump functions means more hyperparameters to learn during training, so in practice, one must maintain a balance between too many hyperparameters and choosing $n_{1}$ and $n_{2}$ to be “large enough.”
Our second test verifies that our kernel performs as well or better than competing kernels within a GP across a variety of ground truth data-generating mechanisms in terms of root mean square error (RMSE) and the continuous rank probability score [CRPS; 49] for out-of-sample test data. This assessment involves first generating data from a mean-zero Gaussian process with a specified kernel. Four kernels are considered: non-sparse and stationary (denoted “scenario” 1 or S1), sparse and stationary (denoted S2), sparse and nonstationary (denoted S3), and non-sparse and nonstationary (denoted S4). See Appendix E for more details and sample draws from each GP. For S1-S4, the synthetic data drawn from the parent distribution represent a ground truth function; we generate $100$ draws, each of which has $50$ (randomly sampled) training measurements and $300$ (regularly spaced) test data points for validation. Each training data set includes the addition of white noise, which is approximately 10% of the signal variance. For each scenario and each synthetic draw, we fit a series of constant-mean GP models to the training data, mirroring the example given in Figure 1: a stationary Matérn kernel (“model” 1 or M1), the sparse general Vecchia (SGV) approximate GP method [32] with a stationary Matérn kernel (M2), our sparsity-discovering kernel proposed from Section 2 with a stationary Matérn core kernel (setting $n_{1}=4$ and $n_{2}=4$; M3), and the same sparsity-discovering kernel with the nonstationary core kernel proposed in Equation B.1 (M4). See Appendix E for further details on training the various Gaussian process models.
Figure 3: Results for the synthetic data tests described in Section D.2. The ground truth “scenarios” are labeled on the $x$-axis. Here, we show the mean and central 90% of continuous rank probability scores [CRPS; 49] and root mean square error (RMSE) across 100 synthetic data sets for three competing methods relative to fitting a Gaussian process with a stationary Matérn kernel (M1). Scores are tallied for each of the data generating mechanisms describe in Table D.1.
In order to ease comparison, we show CRPS and RMSE for M2-M4 relative to corresponding quantities from M1: relative scores less than one indicate that RMSE and CRPS are smaller (and therefore better) compared with a stationary GP. Results are shown in Figure 3, where we show the average (relative) scores and the 5%-95% range over the $100$ draws. When the underlying data are non-sparse and stationary (Scenario 1), i.e., M1 is the “correct” model, methods M2-M4 all perform as well as M1 in terms of both CRPS and RMSE. This makes sense for the Vecchia approximation, since much of the underlying theory is derived for traditional (i.e., non-sparse) stationary kernels; it is reassuring that our kernel (which is nonstationary by definition for both M3 and M4) performs well for stationary data. When the ground truth is sparse and stationary (Scenario 2), M3 and M4 perform similarly–albeit slightly better–than M1, while the Vecchia approximation (M2) performs slightly worse than M1. However, when the underlying true state involves nonstationarity (Scenarios 3 and 4), the nonstationary kernels (M3 and M4) yield a noticeable improvement over a stationary GP, particularly in terms of CRPS. In contrast, M2 is noticeably worse for these scenarios, which suggests that the standard configuration of the sparse general Vecchia approximation is sub-optimal for non-stationary data. The advantage of our kernel is particularly noticeable when the underlying ground truth is nonstationary and sparse.
In summary, these comparisons verify that our kernel performs as well or better than more traditional kernels within a GP across a variety of ground truth data-generating mechanisms. Since each of the data-generating mechanisms considered (different combinations of stationary/nonstationary and sparse/nonsparse) are special cases of our kernel (as described in Section 2.3), this furthermore verifies that our kernel can flexibly learn the true underlying state of a given data set.
4 Application: space-time prediction of daily temperature
Probabilistic machine learning methods like GPs are well-suited for conducting spatio-temporal prediction of in situ measurements of daily maximum temperature. While these data are the gold standard in the Earth sciences, their irregular sampling presents a challenge for climate and atmospheric scientists, who often require data on regular space-time units. To address this need, there are a large number of so-called gridded data products that use (in many cases) naïve statistical methods to conduct spatio-temporal interpolation. For example, the Livneh et al. (2015) [50, 51] data product (henceforth L15) provides daily estimates of surface temperature on a regular grid over the United States from 1950 to present. The L15 data are generated using inverse-distance weighted averaging [52] separately for each day’s measurements before applying a post-hoc correction for orography; no measures of uncertainty quantification are provided. GPs provide many value-added benefits to conducting space-time prediction for temperature measurements relative to the L15 methodology: data-driven (not data-agnostic) weights for spatial averaging; nonstationary weights for interpolation; the ability to account for multiple-day autocorrelation in temperature records; a flexible prior mean function to capture known spatio-temporal structure and systematically estimate relationships between daily temperature and orography; and most importantly uncertainty quantification.
4.1 Spatial modeling of data time slices
We first apply the kernel $C_{z}$ and the Bayesian model to time slices of daily maximum temperature (C) from the GHCN-D database [53], considering stations that lie within the contiguous United States (CONUS). These measurements are the same input data used to construct the L15 data product. We identify a set of $N_{\text{space}}=1174$ GHCN-D stations with high-quality records (see Figure F.1a.) and randomly select 60 days for training: one from each month over 2001-2005. In this case, the input space collapses down to two-dimensional coordinates, longitude and latitude, such that $\mathcal{X}\subset\mathbb{R}^{2}$. For test data, we use measurements from $N_{\text{space, test}}=1444$ records not included in the training records.
We refer the interested reader to Appendix F.1 for full details on the prior mean function and specific kernel used for the time slices. In summary, the prior mean is a linear function of orographic features and distance to the nearest coastlines to account for known relationships between these features and daily temperature. The core kernel is the locally anisotropic nonstationary kernel defined in [45] wherein the signal variance and anisotropy length scales are suitably transformed linear functions of elevation and distance to the coast. For the sparsity-inducing kernel, following the guidance of Section D.1 we err on the side of choosing $n_{1}$ and $n_{2}$ to be larger than we might expect. We consider three configurations, each with 100 total bump functions: (1) $n_{1}=2$ and $n_{2}=50$; (2) $n_{1}=4$ and $n_{2}=25$; and (3) $n_{1}=10$ and $n_{2}=10$. Furthermore, we consider three upper limits on the Wendland truncation radii (5, 10, and 15 units of degrees longitude/latitude) to ensure that the resulting kernel has the chance to impose sparsity on the data. Ultimately, our specification results in five hyperparameters for the prior mean function, twelve hyperparameters for the core kernel, one hyperparameter for the noise variance, and $5033+n_{1}n_{2}(d+3)=3+100\times 5=$ 503 hyperparameters for the sparse kernel: in total there are $5+12+1+503=521$ hyperparameters for the GP applied to each time slice.
Results. Before analyzing all 60 time slices, we first looked at the January 14, 2001 time slice (training and test data shown in Figure 5) to determine an optimal configuration of $n_{1}$ and $n_{2}$ as well as the Wendland length scale upper bound. Three $n_{1}\times n_{2}$ configurations ($10\times 10$, $4\times 25$, $2\times 50$) and three upper bound thresholds result in nine different sparse-inducing GP models. For each model, the MCMC was run for 5,000 iterations, with the first 4,001 discarded as burn-in; convergence of the MCMC was assessed graphically with no issues detected. All subsequent calculations were based on the posterior median of the hyperparameters. The log of the marginal likelihood of the training data, conditioned on the posterior median, the resulting sparsity of the training data covariance, as well as out-of-sample root mean square error (RMSE) and the continuous rank probability score [CRPS; 49] for the test data are shown in Figure 4 for each of the nine models. For comparison, Figure 4 also shows corresponding quantities from a similar Gaussian process model with the same prior mean and core kernel but without the sparse kernel (dashed lines). A few interesting results emerge from the different models considered: first, as expected, the sparsity increases monotonically with the Wendland upper bound; however, unexpectedly, there is a clear trend in the sparsity across the $n_{1}\times n_{2}$ configurations, wherein larger values of $n_{1}$ lead to a smaller fraction of nonzero elements in the covariance. The in-sample log-likelihood and out-of-sample RMSE/CRPS all greatly prefer longer Wendland length-scales (upper bound of 10 or 15), and from a testing standpoint the sparse nonstationary GPs perform as well as the nonsparse nonstationary GP (with about $1/3$ the sparsity). As a compromise between good out-of-sample performance and sparsity, we conclude that the $n_{1}=4$ and $n_{2}=25$ configuration with upper bound 10 is the “optimal” configuration to use for in situ measurements of daily maximum temperature.
[…]
Figure 5: Training and test data for the January 14, 2001 time slice of daily maximum temperature (panels a. and b., respectively). Panels (c) and (d) show the spatial distribution of interpolated daily maximum temperature for both the inverse-distance weighting and our method. The inset text in panels (c) and (d) shows the RMSE for the test data using each method.
Using the optimal configuration, we show visualizations of the prior mean and covariance hyperparameters for the January 14, 2001 time slice in Figure G.2. Since the prior mean function and core kernel parameters are specified as functions of physical, topographical covariates, we can visualize the implied surfaces for the mean function, core kernel signal variance, core kernel anisotropy length-scales, and implied correlation between a set of reference locations and all other points in a $0.25^{\circ}\times 0.25^{\circ}$ longitude-latitude grid over the United States. These quantities (see Figure G.2) provide useful scientific information regarding the first- and second-order properties of daily temperatures, and how these vary according to topography. We can furthermore see how the bump functions impose regions of distance-unrelated correlations across the United States. Finally, we compare predictions for a $0.25^{\circ}\times 0.25^{\circ}$ longitude-latitude grid in Figure 5 for the L15 inverse-distance methodology with corresponding GP-based estimates; see Figure 5(c) and (d). Our methodology yields a clear improvement over L15: both with respect to reduced root mean square error for the test data and also for interpolation of daily maximum temperature in mountainous terrain. For example, note that the inverse-distance weighting completely fails to characterize the topographic variability in, e.g., Nevada and California. Furthermore, it is clear that the L15 method “under-smooths” the data in the eastern United States, particularly along the Gulf Coast.
Now stepping back to consider all 60 time slices, we show corresponding in-sample log-likelihood, sparsity, and out-of-sample metrics in Table 1. For reference, we also summarize RMSE for the L15 method. It is once again clear that a sparse nonstationary GP vastly outperforms L15: on average, the errors are roughly $100(3.00-2.54)/3.00=15.3\%$ smaller (furthermore, the nonstationary GP errors are uniformly smaller than the L15 errors across all time slices). As before, the $n_{1}=4$ and $n_{2}=25$ configuration results in about 37% sparsity, on average.
Table 1: Out-of-sample root mean square error (RMSE) and continuous rank probability score (CRPS) as well as in-sample log-likelihood and sparsity (defined as the fraction of the covariance that is nonzero) aggregated over all 60 time-slices. For comparison, we show RMSE using the inverse-distance weighting methodology of L15. The columns show the average, minimim, and maximum taken over the 60 time-slices.
Method Metric Average Minimum Maximum
L15 RMSE (test data; C) 3.00 2.38 4.22
Nonstationary, sparse GP RMSE (test data; C) 2.54 1.89 4.16
CRPS (test data; no units) 1.50 1.11 2.92
Log-likelihood (training data; no units) -2673.77 -3280.88 -2303.74
Sparsity (fraction of nonzero covariance) 0.37 0.31 0.46
4.2 Spatio-temporal modeling
We next apply the Bayesian model proposed in Eq. 3.1 to spatio-temporal measurements from the $N_{\text{train}}=1174$ stations over 2001-2005 ($N_{\text{time}}=1826$ days), such that the input space is $\mathcal{X}=\mathcal{S}\times\mathcal{T}$ (two-dimensional space and one dimension for time), i.e., $d=3$. To make the calculation more feasible within the allotted time frame, we pick every second data point which yields a total of $N=1,038,723$ non-missing daily measurements for training. We again refer the interested reader to Appendix F.2 for full details on the prior mean function and specific kernel used for the spatio-temporal measurements. In summary, in light of the fact that we are modeling high-dimensional space-time data and our kernel can discover and capitalize upon sparsity, it is to our benefit to propose a prior mean function that explains as much of the known spatial and temporal structure in the data as possible. The most prominent structure in a data set involving temperature is the annual cycle, wherein daily maximum temperatures in the Northern Hemisphere are generally lowest in the winter and highest in the summer. As in Section 4.1, there is an underlying geospatial climatology to measurements of temperature including an implicit zonal structure wherein spatial locations farther from the equator are generally cooler, a strong influence from orographic variability, and the influence of coastlines.
To account for each of these features, we propose a flexible prior mean function using natural cubic splines with spatially-varying coefficients defined by thin-plate spline basis functions and orography. After some trial and error and in order to maintain a compromise between sufficient flexibility and overfitting, we end up with $1899$ hyperparameters for the prior mean: one for every $\approx 550$ data points. Using ordinary least squares, the prior mean function explains 94.1% of the variability in the daily temperature data and removes meaningful autocorrelations beyond $\approx 5$ days (see Figure F.1). In other words, the variability explained by a flexible prior mean function “increases” the sparsity that can be leveraged by the kernel. The space-time core kernel is similar to the one in Equation F.1, albeit generalized to a three-dimensional input space, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})$, with stationarity in the time direction. It is important to note that this specification yields a non-separable space-time kernel, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})\neq C({\bf s},{\bf s}^{\prime})\times C(t,t^{\prime})$. For the sparse kernel, the Wendland component is again anisotropic with coordinate-specific length scales (and coordinate-specific truncation limits); following Section 4.1 we use 10 units for the spatial truncation limits and 4 days for the temporal truncation limits based on the fact that the prior mean function explains autocorrelation over longer time lags. We again use 100 bump functions with $n_{1}=4$ and $n_{2}=25$; however, based on the fact that the prior mean explains so much of the temporal variability, we restrict the bump functions to live in the spatial domain only. This yields $521+1+1=523$ hyperparameters for the kernel; combined with the prior mean hyperparameters, we have $2,422$ hyperparameters to be learned in training with MCMC.
Results. Training the final GP involved 1000 MCMC iterations, which translated into approximately 7000 likelihood evaluations on 256 NERSC Perlmutter A100 GPUs. Although the time cost varies slightly based on the current hyperparameters and the induced sparsity structure of the covariance matrix, generally speaking, a representative likelihood took 66 seconds to compute: 37 seconds for distributed covariance computation, about 9 seconds for the MINRES (Minimal Residual Method, [54]) linear solve, 2 seconds for the log-determinant calculation, with the rest of the time spent on mean and noise function calculations and necessary transformations between formats. The observed sparsity was approximately 0.0006 (nonzero elements/matrix size). In total, this led to an approximate computation time of 128 hours – well in line with the training of large neural networks or large language models.
Figure 6: Test data with posterior predictions and uncertainties for seven selected selected sites over 2003-2005 (2001-2002 omitted for visibility of results). The shaded band represents the three-sigma uncertainty interval.
To summarize the results of this analysis, we selected seven test sites from across the United States and assessed the time-evolving behavior of daily maximum temperatures over 2001-2005; see Figure 6 (the first two years are omitted for visibility of results). The training sites with test locations are shown on the top left, with the test data and posterior predictions shown in each thumbnail. The most obvious feature of the predictions is the annual cycle, which is well-characterized by our prior mean function. However, note that the specific shape of the annual cycle is flexibly modeled across sites: temperatures in Houston, TX (top right) increase sharply in late winter and decrease sharply in late fall, while temperatures in Portland, OR (middle left) increase and decrease much more slowly between the cold and warm season. The GP predictions maintain the large day-to-day variations present in daily maximum temperatures, which is reflected in the data. Also note that the borrowing of strength over space and time allows the GP to learn the time-varying structure even when there are large chunks of missing data in the traning data, e.g., in the first half of 2003 at the Houston, TX gauged location. Finally, the three-sigma GP uncertainties, represented by the red shaded band in Figure 6, are clearly nonstationary as well: note the larger magnitude of uncertainties at, e.g., Houston, TX compared to the smaller uncertainties for the Berkeley, CA site.
5 Discussion
In this work, we have proposed a highly flexible kernel that encodes and discovers both sparsity and nonstationarity. We demonstrate how the proposed kernel can be embedded within a Bayesian Gaussian process stochastic model, as well as flexible MCMC tools for GP training. Our methodology outperforms a variety of competing methods across different ground truth data and furthermore provides better prediction errors than state-of-the-art methods in the Earth sciences for measurements of daily maximum temperature. Our application of an exact GP to one million data points is on par with the largest GPs applied to date. In the scope of exact GPs, [55] scaled to 1.3 million data points by taking advantage of GPU-accelerated matrix-vector products in Conjugate-Gradient solver for faster training. For this method to work, the covariance matrix has to be well-conditioned which is only the case for some stationary kernels. In addition, their method has a natural maximum scaling due to the sheer size of the dense covariance matrix. Our proposed method, in contrast, takes advantage of naturally occurring sparsity, has shown favorable weak and strong scaling properties [40], and works in harmony with non-stationary kernels.
Depending on the level of sparsity that is discovered by the proposed kernel, computational challenges arise during the log-determinant calculation and the linear system solution. The log-determinant calculation is currently performed using a Lanczos quadrature algorithm [56]. Although this estimation is fast, it is approximate, which may or may not affect the number of MCMC iterations and the solution; however, among methods to train a GP, the MCMC will likely handle this inaccuracy most robustly. Ideally, one might deploy a sparse Cholesky or LU decomposition, but current algorithms produce (LU and Cholesky) factors that are not sparse enough to be stored. We are working on a Python LU multi-node decomposition that would allow fast and exact computation of the log-determinant. LU or Cholesky decomposition would also allow for a fast and exact linear system solution, which is currently accomplished by the MINRES algorithm. Given our kernel and covariance structure, the conjugate-gradients method converges poorly, and common preconditioners are more time-consuming than MINRES, in our case.
With regards to scalability beyond one million data points, [40] derived a formula to predict how much time it would take to calculate the covariance matrix for $N$ data points: $T=\frac{N^{2}t_{b}}{2nb^{2}}$. Here, $N$ is the size of the dataset, $t_{b}$ is the time it takes for a covariance batch to be computed, $n$ is the number of workers, and $b$ is the batch size. The validity of this formula has been well-tested in our experiments. Assuming $n=20,000$ available GPUs that can calculate a batch of size $b=15,000$ in 2 seconds, we could calculate the covariance matrix for 100 million data points in $2$ seconds; with the current algorithm, the log-determinant and the MINRES could be calculated in about 1000 seconds each. This results in a log marginal likelihood evaluation time of just over an hour. Whether or not this is still in the realm of feasibility depends on the total available computing time. If the number of hyperparameters is relatively small, the MCMC might converge after around 500 iterations, which amounts to just over 3 weeks of total computing time. This toy example reveals that a 100 million-data-point GP is on the horizon. Note, however, that this calculation assumes constant sparsity similar to the one observed in the real data application in Section 4.2. Early evidence suggests that larger datasets will allow the proposed kernel to discover more sparsity. While this will not affect the calculation of the covariance matrix, it vastly reduces linear system solutions and log-determinant calculations.
Declarations
Funding. This research was supported by three programs:
  1. 1.
    The Director, Office of Science, Office of Biological and Environmental Research of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 and by the Regional and Global Model Analysis Program, and the Calibrated and Systematic Characterization, Attribution, and Detection of Extremes (CASCADE) Scientific Focus Area, within the Earth and Environmental Systems Modeling Program.
  2. 2.
    The Center for Advanced Mathematics for Energy Research Applications (CAMERA), which is jointly funded by the Advanced Scientific Computing Research (ASCR) and Basic Energy Sciences (BES) within the Department of Energy’s Office of Science, under Contract No. DE-AC02-05CH11231.
  3. 3.
    This research was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory.
The research also used resources of the National Energy Research Scientific Computing Center (NERSC), also supported by the Office of Science of the U.S. Department of Energy, under Contract No. DE-AC02-05CH11231.
Conflict of interest/Competing interests. The authors declare no competing financial interests.
Ethics approval and consent to participate. Not applicable. Consent to participate. Not applicable. Consent for publication. Not applicable.
Availability of data and materials. The in situ temperature records supporting this article are based on publicly available measurements from the National Centers for Environmental Information (ftp://ftp.ncdc.noaa.gov/pub/data/ghcn/daily/).
Code availability. All data analysis in this manuscript was conducted using open-source programming languages and software (namely, R and Python). The BayesNSGP software package for R is available at https://github.com/danielturek/BayesNSGP. The gpCAM Python package is available via “pip install gpcam” or at https://github.com/lbl-camera/gpCAM.
Authors’ contributions. MDR contributed initial conception of the study and led all analyses, figure generation, and writing. MMN contributed to the initial conception, developed the Python algorithm for scalable GPs, ran all models on the Perlmutter supercomputer, and added various content to the manuscript. HL contributed to discussion of the methodological development and improvement of mixing, and writing. RP contributed by optimizing the performance of the scalable-GP code and streamlined the code deployment on Github and PyPI. All authors contributed to the interpretation of study results and editing of the study’s writing and figures.
Appendix A Proof of Proposition 1
Proposition 1. The kernel $C_{y}$ has $w=\min\{w_{0},w_{\text{core}}\}$ continuous derivatives at zero, where $f_{0}$ has $w_{0}$ continuous derivatives at the origin and $C_{\text{core}}$ has $w_{\text{core}}$ continuous derivatives at zero.
Proof.
The proof follows directly from the general Leibniz rule, which generalizes the product rule for higher-order differentials. It is well-established and proven by induction. The general Leibniz rule states
$\partial^{l}(gh)=\sum_{k}^{l}\binom{l}{k}(\partial^{k}g)(\partial^{l-k}h)$
for two functions $g$ and $h$. If we assume the $g$ and $h$ are $n$ and $m$ times differentiable respectively, the sum only exists for $l\leq\min\{m,n\}$. This can be shown by considering three scenarios:
$l\leq\min\{m,n\}$ (A.1)
$l>\min\{m,n\}\land l<\max\{m,n\}$ (A.2)
$l\geq\max\{m,n\}.$ (A.3)
Scenario (A.2) and (A.3) both lead to terms $\partial^{k}g$ with $k>n$ or $\partial^{l-k}h$ with $l-k>m$. All terms exist and are continuous only for scenario (A.1), which concludes the proof. ∎
Appendix B Parametric nonstationary kernel
In the absence of specific knowledge regarding the form of the core kernel, a natural and flexible choice that performs well in practice is the so-called “parametric” nonstationary kernel [46] derived in [15] and later extended in [45]. This ultra-flexible kernel allows all aspects of the kernel to vary according to positions in the input data:
$C_{\text{core}}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{\text{core}})=\sigma({\bf x})\sigma({\bf x}^{\prime})\frac{\big{(}\Sigma({\bf x})\Sigma({\bf x}^{\prime})\big{)}^{d/4}}{\left(\frac{\Sigma({\bf x}^{\prime})+\Sigma({\bf x}^{\prime})}{2}\right)^{d/2}}\mathcal{M}_{\nu}\left(\sqrt{Q({\bf x},{\bf x}^{\prime})}\right),\hskip 12.91663pt{\bf x},{\bf x}^{\prime}\in\mathcal{X},$ (B.1)
where
$Q({\bf x},{\bf x}^{\prime})=\frac{({\bf x}-{\bf x}^{\prime})^{\top}({\bf x}-{\bf x}^{\prime})}{\frac{1}{2}[\Sigma({\bf x})+\Sigma({\bf x}^{\prime})]}=\frac{||{\bf x}-{\bf x}^{\prime}||^{2}}{\frac{1}{2}[\Sigma({\bf x})+\Sigma({\bf x}^{\prime})]},$
and $\mathcal{M}_{\nu}(\cdot)$ is the Matérn correlation function with smoothness $\nu$. In Eq. B.1, $\sigma(\cdot)$ is the signal standard deviation, and $\Sigma(\cdot)$ controls the length-scale of correlation. This kernel is locally isotropic; note that it is straightforward to generalize Eq. B.1 to be locally anisotropic [see, e.g., 45]. One can then parameterize $\sigma(\cdot)$ and $\Sigma(\cdot)$ to be log-linear functions of a set of basis functions, e.g.,
$\log\sigma({\bf x})=\log\sigma_{0}+\sum_{m=1}^{M_{\sigma}}s_{m}({\bf x})\phi_{\sigma}^{m},\hskip 21.52771pt\log\Sigma({\bf x})=\log\Sigma_{0}+\sum_{m=1}^{M_{\Sigma}}s_{m}({\bf x})\phi_{\Sigma}^{m},$
where $\sigma_{0}$ and $\Sigma_{0}$ are the baseline signal standard deviation and length-scale, respectively, $\{s_{m}(\cdot):m=1,\dots,M\}$ is a set of generic basis functions (polynomials, splines, etc.) and $\{\phi_{(\cdot)}^{m}:m=1,\dots,M_{(\cdot)}\}$ is a set of basis function coefficients. Of course, the basis functions for $\sigma(\cdot)$ and $\Sigma(\cdot)$ could be different as well. Using this framework, the hyperparameters for the core kernel are then ${\boldsymbol{\theta}}_{\text{core}}=(\sigma_{0},\Sigma_{0},\{\phi_{\sigma}^{m}\},\{\phi_{\Sigma}^{m}\})$.
However, when using the generic nonstationary core kernel defined in Equation B.1 some care is required to regularize the hyperparameters to guard against overfitting. Even if $M_{\sigma}$ and $M_{\Sigma}$ are not too large, if the “true” signal standard deviation and length scale are stationary this could result in overfitting in the case of limited data. To this end, we apply a simple regularization prior to the $\{\phi_{(\cdot)}^{m}\}$:
$\phi_{(\cdot)}^{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,v_{(\cdot)})\hskip 12.91663pt\text{where}\hskip 12.91663ptv_{(\cdot)}\sim U(0,10^{5}).$
This prior distribution is equivalent to applying an $L_{2}$ penalty on the $\{\phi_{(\cdot)}^{m}\}$ as is used in ridge regression in a Frequentist setting; in a Bayesian setting, $v_{(\cdot)}$ is inferred from the data (i.e., it is updated within the MCMC). This prior has the effect of regularizing the basis function coefficients: when $v_{(\cdot)}$ is small, the $\phi_{(\cdot)}^{m}$ are “shrunk” towards zero; on the other hand, when $v_{(\cdot)}$ is large the coefficients are essentially unrestricted.
Appendix C Gaussian process training via Markov chain Monte Carlo
Generally speaking, the posterior distribution (5) is not available in closed form regardless of prior choice, and so we must resort to Markov chain Monte Carlo (MCMC) methods to conduct inference on ${\boldsymbol{\beta}}$ and ${\boldsymbol{\theta}}$ [47]. In practice, the specific Markov chain Monte Carlo algorithm used for any given data set will be customized to find a suitable balance between computational time and efficiency (or “mixing”) of the chain; however, we found that the following algorithm works well. First, adaptive block Metropolis-Hastings random walk samplers [57, 58] are used on the following groups of hyperparameters:
  1. 1.
    The prior mean function coefficients ${\boldsymbol{\beta}}$ (note that a closed-form Gibbs update is available for ${\boldsymbol{\beta}}$ under the likelihood and prior choice described above; we found that an adaptive Metropolis update worked just as well and obviated additional computational steps).
  2. 2.
    The core kernel hyperparameters ${\boldsymbol{\theta}}_{\text{core}}$, the error variance hyperparameters ${\boldsymbol{\theta}}_{z}$, and Wendland hyperparameters $s_{0}$ and $r_{0}$.
  3. 3.
    The bump function positions $\{{\bf h}_{ij}\}$ and radii $\{r_{ij}\}$.
A fourth block updates all of the amplitudes $\{a_{ij}\}$ in a single step. Since these are binary variables, the proposal distribution matches the prior, i.e., a proposed $a_{ij}^{*}$ is drawn from a Bernoulli distribution with success probability $\pi_{ij}^{\text{curr}}$ (the current value of the prior probabilities). Lastly, we can use closed-form Gibbs updates for the prior probabilities $\{\pi_{ij}\}$ because their full conditional distribution can be derived in closed form. Specifically, the distribution of a single $\pi_{ij}$ conditioned on all other hyperparameters and the data is
$p(\pi_{ij}|{\bf z},{\boldsymbol{\beta}},{\boldsymbol{\theta}}_{-\{\pi_{ij}\}})\propto p({\bf z}|{\boldsymbol{\beta}},{\boldsymbol{\theta}})p({\boldsymbol{\beta}},{\boldsymbol{\theta}})\propto p(\pi_{ij})p(a_{ij}|\pi_{ij})$
(here, proportionality refers to quantities involving $\pi_{ij}$). Noting that the prior for $\pi_{ij}$ (uniform on the unit interval) can be re-written as a Beta(1,1)11(1,1)( 1 , 1 ) distribution, i.e.,
$p(\pi_{ij})=\frac{\Gamma(1+1)}{\Gamma(1)\Gamma(1)}\pi_{ij}^{(1-1)}(1-\pi_{ij})^{(1-1)},\hskip 21.52771pt0<\pi_{ij}<1,$
and that the Bernoulli distribution $p(a_{ij}|\pi_{ij})$ is
$p(a_{ij}|\pi_{ij})=\pi_{ij}^{a_{ij}}(1-\pi_{ij})^{(1-a_{ij})},$
as a function of $\pi_{ij}$ the product becomes
$\begin{array}[]{rcl}p(\pi_{ij})p(a_{ij}|\pi_{ij})&\propto&\pi_{ij}^{(1-1)}(1-\pi_{ij})^{(1-1)}\pi_{ij}^{a_{ij}}(1-\pi_{ij})^{(1-a_{ij})}\\[3.65973pt] &\propto&\pi_{ij}^{1-1+a_{ij}}(1-\pi_{ij})^{1-1+1-a_{ij}}\\[3.65973pt] &\propto&\pi_{ij}^{(1+a_{ij})-1}(1-\pi_{ij})^{(2-a_{ij})-1}\end{array}$
which is proportional to a Beta$(1+a_{ij},2-a_{ij})$ distribution. Hence, the prior probabilities can be updated by sampling from a Beta$(1+a_{ij}^{\text{curr}},2-a_{ij}^{\text{curr}})$ distribution.
Appendix D Synthetic data examples
We now describe two synthetic data examples to demonstrate the performance of our proposed kernel under mis-specification of the number of bump functions (Section D.1) and relative to competing methods across a variety of ground-truth data states (Section D.2). In Section D.2, we assess performance quality using two metrics: (1) root mean square error (RMSE) for the test set, which assesses the pointwise error in posterior mean predictions; and (2) the continuous rank probability score [CRPS; 49] for the test set, which (unlike the RMSE) assesses both precision and sharpness of predictions. For example, the CRPS rewards a bad prediction with large uncertainty over a bad prediction with small uncertainty. Note that the reported CRPS scores are calculated for each test data point and then averaged across all test data.
D.1 Mis-specified bump functions
First, we conduct a test to assess the performance of our proposed kernel when the number of bump functions is mis-specified. For a real-world data application, one of course does not know how many bump functions to specify, which is a requisite a priori choice that must be made when using our methodology. The goal of this subsection is to provide guidance for making this decision.
Figure D.1: Results for the synthetic data test described in Sections D.1. The ground truth function and noisy measurements are shown in panel (a); panels (b) and (c) show best estimates of the bump functions (posterior medians) and posterior predictions with uncertainty (posterior mean and 90% credible intervals), respectively. Results are shown when the kernel has too few bump functions (left), the correct number of bump functions (center), and too many bump functions (right).
The synthetic data test proceeds as follows. First, we suppose the ground truth function is a piecewise-constant function on the unit interval, defined as
$y(x)=\left\{\begin{array}[]{cc}-1&\text{if }x\leq 0.25\text{ or }0.5<x\leq 0.75\\ 1&\text{if }0.25<x\leq 0.5\text{ or }x>0.75,\end{array}\right.$
and we have noisy measurements from this function for training a GP (see Figure D.1a.). The form of this ground truth function suggests that there should be two sets of two bump functions: $y(x)$ is perfectly correlated in $(0,0.25]$ and $(0.5,0.75]$; furthermore, $y(x)$ is perfectly correlated in $(0.25,0.5]$ and $(0.75,1)$. In other words, $n_{1}=2$ and $n_{2}=2$. We then fit our proposed kernel to the noisy data assuming we know nothing about its ground truth, under three configurations of bump functions: too few bump functions, where we set $n_{1}=1$ and $n_{2}=2$; the right number of bump functions, using $n_{1}=2$ and $n_{2}=2$; too many bump functions, using $n_{1}=4$ and $n_{2}=4$. Notice that although the second configuration is the only one with the correct number of bump functions, in all cases we infer the bump functions characteristics (positions and radii) directly from the data. When fitting, we suppose the core kernel is a stationary Matérn with smoothness $\nu=2.5$.
Results are shown in Figure D.1(b) and (c), where we show both the posterior median estimates of the bump functions and the posterior predictions of $y(\cdot)$ with uncertainty, respectively. First, consider the bump function estimates in Figure D.1(b). When there are too few bump functions, the training correctly identifies the position and size of the bump functions for part of the domain, i.e., that $(0.25,0.5]$ is correlated with $(0.75,1)$, while the kernel does not have the flexibility to capture the distance-unrelated correlations in $(0,0.25]$ and $(0.5,0.75]$. As such, the predictions capture the ground truth function with very low uncertainty in $(0.25,0.5]$ and $(0.75,1)$ but revert to the sparse stationary kernel in the other parts of the domain – effectively overfitting the data in $(0,0.25]$ and $(0.5,0.75]$. When there are the correct number of bump functions, the training correctly identifies where these bump functions should go and how large they should be; consequently, posterior predictions are correctly calibrated over the entire unit interval. Finally, when there are too many bump functions, the algorithm correctly determines both the size and position of the bump functions, but also that only four of the 16 total bump functions should be activated. The other 12 are appropriately zeroed out. Posterior predictions are effectively identical to those generated using the correct number of bump functions.
In summary, this example verifies two important points: (1) our proposed sparse kernel can both identify where and how big the bump functions should be, and (2) the regularization method for the bump functions described in Section 3.1 effectively eliminates the effect of any bump functions that may be unnecessary. In other words, for a real data example, we recommend specifying more bump functions than one might think is truly necessary (i.e., err on the side of making $n_{1}$ and $n_{2}$ large). Of course, more bump functions means more hyperparameters to learn during training, so in practice one must maintain a balance between too many hyperparameters and choosing $n_{1}$ and $n_{2}$ to be “large enough.”
D.2 Comparison with competing methods
Next, we generate data from a variety of underlying data generating mechanisms and compare our approach with competing methods. The data generating mechanisms are designed to span different combinations of sparsity and nonstationarity that may be present in real-world data sets. Specifically, we generate data from mean-zero multivariate Gaussian distributions where the “true” covariance is defined as follows:
  1. S1
    Non-sparse and stationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma^{2}\mathcal{M}_{\nu}(||{\bf x}-{\bf x}^{\prime}||)$, where $\mathcal{M}_{\nu}(\cdot)$ is the Matérn kernel with smoothness $\nu$.
  2. S2
    Sparse and stationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma^{2}f_{0}({\bf x},{\bf x}^{\prime};r_{0})$, where $f_{0}$ is as in Eq. 2.1.
  3. S3
    Sparse and nonstationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma({\bf x})\sigma({\bf x}^{\prime})f_{0}({\bf x},{\bf x}^{\prime};r_{0})$, where the signal variance $\sigma^{2}(\cdot)$ depends on the input locations.
  4. S4
    Non-sparse and nonstationary: here we take $C_{y}$ to be the nonstationary kernel described in [45].
Here we use S1-S4 to denote different “scenarios”; see Appendix E for more details on the data generating mechanisms and specific hyperparameters used. Sample draws from the true parent distribution for a one-dimensional input space are shown in Supplemental Figure E.1. In each case, the synthetic data drawn from the parent distribution represents a ground truth function; we can therefore generate $N_{\text{train}}=50$ (randomly sampled) and $N_{\text{test}}=300$ (regularly spaced) data points for validation. Each training data set will include the addition of white noise that is approximately 10% of the signal variance. For each scenario, we generate $N_{\text{rep}}=100$ replicates or draws from the ground truth distribution; note that the locations of the training data are different across each replicate.
For each scenario and each replicate, we fit a series of constant-mean GP models to the synthetic data, mirroring the example given in Figure 1 (the “M” denotes a statistical model):
  • M1
    Stationary Gaussian process with a Matérn kernel
  • M2
    Sparse general Vecchia (SGV) approximate GP method as proposed in [32]
  • M3
    The sparsity-discovering kernel proposed in Section 2, using a stationary Matérn for the core kernel
  • M4
    The sparsity-discovering kernel proposed in Section 2, using the nonstationary core kernel defined in Equation B.1 where the $s_{m}(\cdot)$ are natural smoothing splines [59] and $M_{\sigma}=M_{\Sigma}=4$.
We note that M1 and M2 could in principle involve nonstationary kernels; however, we maintain stationary kernels for two reasons. First, M1 is primarily used here as a reference for what one can obtain with a “default” (stationary and exact) GP when working with smaller data sets; the whole point of this methodology is that M1 will become unusable for much larger data sets. Second, all of the underlying theory for Vecchia approximations [32] has been derived for stationary kernels. It is, of course, possible to use the Vecchia framework with a nonstationary kernel [see, e.g., 45]; however, without further theoretical developments there are no guarantees that the subjective choices one must make for the Vecchia approximation (size of conditioning sets, ordering of points, selection of nearest neighbors) will result in an appropriate approximation for a nonstationary kernel. The different data-generating mechanisms and statistical models are summarized in Table D.1. Our implementation of M1 and M2 involve standard, off-the-shelf implementations of these methods as enabled by the BayesNSGP software package for R [60]. For M3 and M4, the main user-specified feature is determining how many bump functions are allowed: here, we set $n_{1}=4$ and $n_{2}=4$). In order to ease comparison, we show CRPS and RMSE for M2-M4 relative to corresponding quantities from M1: relative scores less than one indicate RMSE and CRPS are smaller (and therefore better) for the model of interest versus a stationary GP.
Table D.1: Summary of five data-generating mechanisms (or “scenarios”) used to generate synthetic data and four kernels used within a Gaussian process to fit each of $N_{\text{rep}}=100$ repeated draws from each scenario.
Data generating mechanisms Kernels within the Gaussian process
Label Description Label Description
S1 Non-sparse, stationary M1 Stationary Matérn kernel
S2 Sparse, stationary M2 Vecchia approximation with Matérn
S3 Sparse, nonstationary M3 Sparsity-discovering kernel with stationary core
S4 Non-sparse, nonstationary M4 Sparsity-discovering kernel with nonstationary core
Appendix E Synthetic data examples: hyperparameters and sample data sets
Here, we provide specific details on the data generating mechanisms S1-S4 used in Section D.2, including hyperparameter configurations and sample draws from each parent distribution. The input space is $\mathcal{X}=[0,10]$, i.e., $d=1$ and elements of the input space are scalars.
S1: Non-sparse and stationary.
Here, recall that $C_{y}(x,x^{\prime})=\sigma^{2}\mathcal{M}_{\nu}(|x-x^{\prime}|)$. We specify $\sigma^{2}=1$, $\nu=2.5$, and the Matérn length-scale parameter is $\rho=0.5$. The training data include additive Gaussian white noise with variance $\tau^{2}=0.1$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
S2: Sparse and stationary.
Here, $C_{y}(x,x^{\prime})=\sigma^{2}f_{0}(x,x^{\prime};r_{0})$, where $f_{0}$ is as in Eq. 2.1. We specify $\sigma^{2}=1$ and $r_{0}=1.5$. The training data include additive Gaussian white noise with variance $\tau^{2}=0.1$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
S3: Sparse and nonstationary.
Here, $C_{y}(x,x^{\prime})=\sigma(x)\sigma(x^{\prime})f_{0}(x,x^{\prime};r_{0})$, where the signal variance $\sigma^{2}(\cdot)$ depends on the input locations, $f_{0}$ is as in Eq. 2.1, and $r_{0}=0.75$. Specifically,
$\sigma^{2}(x)=0.05(x-5)^{4}+0.001.$
The training data include additive Gaussian white noise with variance $\tau^{2}=0.475$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
S4: Non-sparse and nonstationary.
Here, we use
$C_{y}(x,x^{\prime})=\sigma(x)\sigma(x^{\prime})\frac{\big{(}\Sigma(x)\Sigma(x^{\prime})\big{)}^{1/4}}{\left(\frac{\Sigma(x^{\prime})+\Sigma(x^{\prime})}{2}\right)^{1/2}}\mathcal{M}_{\nu}\left(\sqrt{\frac{|x-x^{\prime}|^{2}}{\frac{1}{2}[\Sigma(x)+\Sigma(x^{\prime})]}}\right),$
where $\nu=2.5$,
$\sigma^{2}(x)=0.2\left(\frac{x-10}{3}\right)^{4}+0.1,$
and
$\Sigma(x)=0.06\left({x}/{3}\right)^{3}+0.03.$
The training data include additive Gaussian white noise with variance $\tau^{2}=0.506$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
Figure E.1: Four sample draws from each of the four parent distributions described in Table D.1.
Appendix F Prior mean and kernel for temperature data
F.1 Spatial modeling of data time slices
Prior mean function for time slices. For time-slices of daily maximum temperature measurements, there are two primary features to account for in the prior mean: relationships with orography and distance from the coast. Specifically, we generally expect (1) inverse relationships between temperature and elevation (via so-called “lapse-rate” theory); and (2) direct relationships between temperature and distance to coastlines. We can furthermore include additional topographic variables of slope and aspect, which respectively describe the direction in which each location faces (ranging from zero to $2\pi$) and the steepness. Collectively, slope and aspect account for the effect of solar radiation and exposure on daily maximum temperatures. The elevation is obtained from the 800m digital elevation map used to generate the PRISM data product; slope and aspect are calculated using the raster package for R [61]. The distance-to-coast (km) variable is calculated by NASA’s Ocean Biology Processing Group [62] at a global grid of $1/16^{\circ}$ using the Generic Mapping Tools package [63] (note that this variable only calculates distances to ocean coastlines and not, e.g, other landlocked bodies of water). Together, the prior mean function is a linear function of these covariates:
$\mu({\bf s})=\alpha_{0}+\alpha_{1}\text{Elev}({\bf s})+\alpha_{2}\text{Slope}({\bf s})+\alpha_{3}\text{Aspect}({\bf s})+\alpha_{4}\text{Dist2Coast}({\bf s}),$
where we update the coefficients $\{\alpha_{j}:j=0,\dots,4\}$ within the MCMC.
Kernel. For the core kernel, we use the locally anisotropic nonstationary kernel defined in [45] wherein the signal variance and anisotropy length scales are suitably transformed linear functions of elevation and distance to the coast. Specifically, for ${\bf s},{\bf s}^{\prime}\in\mathcal{S}$,
$C_{\text{core}}({\bf s},{\bf s}^{\prime};{\boldsymbol{\theta}}_{\text{core}})=\sigma({\bf s})\sigma({\bf s}^{\prime})\frac{\left|{\bf\Sigma}({\bf s})\right|^{1/4}\left|{\bf\Sigma}({\bf s}^{\prime})\right|^{1/4}}{\left|\frac{{\bf\Sigma}({\bf s}^{\prime})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right|^{1/2}}\mathcal{M}_{1.5}\left(\sqrt{({\bf s}-{\bf s}^{\prime})^{\top}\left(\frac{{\bf\Sigma}({\bf s})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right)^{-1}({\bf s}-{\bf s}^{\prime})}\right),$ (F.1)
where $\mathcal{M}_{1.5}$ is the Matérn kernel with smoothness equal to 1.5. We parameterize the spatially-varying signal variance as
$\log\sigma({\bf s})=\phi_{0}+\phi_{1}\text{Elev}({\bf s})+\phi_{2}\text{Dist2Coast}({\bf s}).$
For the anisotropy, we use the eigendecomposition
${\bf\Sigma}({\bf s})={\bf\Gamma}({\bf s}){\bf\Lambda}({\bf s}){\bf\Gamma}({\bf s})^{\top},\hskip 12.91663pt\text{where}\hskip 12.91663pt{\bf\Lambda}({\bf s})=\left[\begin{array}[]{cc}\lambda_{1}({\bf s})&0\\ 0&\lambda_{2}({\bf s})\end{array}\right],{\bf\Gamma}({\bf s})=\left[\begin{array}[]{cc}\cos\gamma({\bf s})&-\sin\gamma({\bf s})\\ \sin\gamma({\bf s})&\cos\gamma({\bf s})\end{array}\right],$
and
$\begin{array}[]{c}\log\lambda_{1}({\bf s})=\omega^{\lambda_{1}}_{0}+\omega^{\lambda_{1}}_{1}\text{Elev}({\bf s})+\omega^{\lambda_{1}}_{2}\text{Dist2Coast}({\bf s}),\\[2.15277pt] \log\lambda_{2}({\bf s})=\omega^{\lambda_{2}}_{0}+\omega^{\lambda_{2}}_{1}\text{Elev}({\bf s})+\omega^{\lambda_{2}}_{2}\text{Dist2Coast}({\bf s}),\\[2.15277pt] \log\frac{\frac{2}{\pi}\gamma({\bf s})}{1-\frac{2}{\pi}\gamma({\bf s})}=\omega^{\gamma}_{0}+\omega^{\gamma}_{1}\text{Elev}({\bf s})+\omega^{\gamma}_{2}\text{Dist2Coast}({\bf s}).\end{array}$ (F.2)
These specific transformations are chosen so that the coefficients $\{\omega_{j}^{(\cdot)}\}$ can be unrestricted and still yield valid ${\bf\Sigma}(\cdot)$. For the sparsity-inducing kernel, the primary choice is to specify $n_{1}$ and $n_{2}$, the number of sets ($n_{1}$) of correlated yet distance-unrelated regions ($n_{2}$). Following the guidance of Section D.1, we err on the side of choosing these to be larger than we might expect. We consider three configurations, each with 100 total bump functions: (1) $n_{1}=2$ and $n_{2}=50$; (2) $n_{1}=4$ and $n_{2}=25$; and (3) $n_{1}=10$ and $n_{2}=10$. Note that in this implementation, for the sparse kernel, we use an anisotropic version of the Wendland kernel that has separate $x$- and $y$-coordinate truncation radii. Furthermore, we consider three upper limits on the Wendland truncation radii (5, 10, and 15 units of degrees longitude/latitude) since we found that otherwise the truncation radii tend to increase without bound in the training. Limiting these radii to be relatively small ensures that the resulting kernel has the chance to impose sparsity on the data.
In summary, we have five hyperparameters for the prior mean function ($\{\alpha_{j}:j=0,\dots,4\}$), twelve hyperparameters for the core kernel (${\boldsymbol{\theta}}_{\text{core}}=\{\phi_{j},\omega_{j}^{\lambda_{1}},\omega_{j}^{\lambda_{2}},\omega_{j}^{\gamma}:j=0,1,2\}$), one hyperparameter for the noise variance. In the sparse kernel, 100 bump functions means that we have $5033+n_{1}n_{2}(d+3)=3+100\times 5=$ 503 hyperparameters. Overall, this yields $5+12+1+503=521$ hyperparameters for the GP applied to each time slice.
F.2 Spatio-temporal modeling
Prior mean function in space and time. In light of the fact that we are modeling high-dimensional space-time data and our kernel can discover and capitalize upon sparsity, it is to our benefit to propose a prior mean function that explains as much of the known spatial and temporal structure in the data as possible. The most prominent structure in a data set involving temperature is the annual cycle, wherein daily maximum temperatures in the Northern Hemisphere are generally lowest in the winter and highest in the summer. As in Section 4.1, there is an underlying geospatial climatology to measurements of temperature including an implicit zonal structure wherein spatial locations farther from the equator are generally cooler, a strong influence from orographic variability, and the influence of coastlines. To account for each of these features, we propose the following prior mean function for each ${\bf x}=({\bf s},t)\in\mathcal{X}=\mathcal{S}\times\mathcal{T}$:
$\mu({\bf x})=\alpha_{0}({\bf s})+\sum_{j=1}^{J}\alpha_{j}({\bf s})v_{j}(t).$ (F.3)
Here, the $v_{j}(\cdot)$ are natural cubic splines that map day $t$ to its Julian day to model the seasonal cycle, i.e.,
$v_{j}(t)\equiv\left\{\begin{array}[]{ll}v_{j}\left(\cos\left([t\mod 365]\frac{2\pi}{365}\right)\right)&j=1,\dots,J/2\\ v_{j}\left(\sin\left([t\mod 365]\frac{2\pi}{365}\right)\right)&j=J/2+1,\dots,J\end{array}\right.$
(swapping in 366 for 365 in leap years). Using sine- and cosine-based splines allows us to flexibly model the seasonal cycle such that the cycle in continuous in $t$; even a relatively low-dimensional representation (i.e., $J\approx 6$) explains roughly 75% of the variability in the daily data. Each of these cubic splines in time are assigned a spatially-varying coefficient, $\{\alpha_{j}(\cdot):j=0,\dots,J\}$, which we model using thin plate splines in the geographic coordinates, elevation, and distance-to-coast:
$\alpha_{j}({\bf s})=\sum_{m=1}^{M_{s}}\delta_{jm}^{s}q^{s}_{m}({\bf s})+\sum_{m=1}^{M_{e}}\delta_{jm}^{e}q^{e}_{m}(e)+\sum_{m=1}^{M_{c}}\delta_{jm}^{d}q^{c}_{m}(c).$ (F.4)
Here, $\{q^{s}_{m}(\cdot)\}$, $\{q^{e}_{m}(\cdot)\}$, and $\{q^{c}_{m}(\cdot)\}$ are sets of thin plate spline basis functions that define nonlinear mappings from the input coordinates ${\bf s}$, elevation $e\equiv e({\bf s})$, and distance-to-coast $c=c({\bf s})$, respectively. The thin plate spline basis functions are considered fixed but their coefficients $\{\delta^{s}_{jm}\}$, $\{\delta^{e}_{jm}\}$, and $\{\delta^{c}_{jm}\}$ are considered hyperparameters and updated in the MCMC.
The linearity assumptions of Equations F.3 and F.4 allow us to write the mean function $\mu({\bf x})$ as a linear function of a set of covariates, as denoted in Section 3.1. Let ${\bf W}_{\text{time}}$ denote a $N_{\text{time}}\times(J+1)$ matrix of cubic spline functions ($J$ of which correspond to the cubic splines and 1 for an intercept), where $N_{\text{time}}=|\mathcal{T}|$, and let ${\bf W}_{\text{space}}$ denote a $N_{\text{space}}\times(1+M_{s}+M_{e}+M_{c})$ matrix of thin plate spline functions with an intercept, where $N_{\text{space}}=|\mathcal{S}|$. The design matrix ${\bf W}$ can then be written as
${\bf W}={\bf W}_{\text{space}}\otimes{\bf W}_{\text{time}}$ (F.5)
(where $\otimes$ denotes the Kronecker product), which is applicable when the observational vector ${\bf z}$ concatenates all daily measurements from each station in turn, i.e.,
${\bf z}=\big{(}z({\bf s}_{1},t_{1}),z({\bf s}_{1},t_{2}),\dots,z({\bf s}_{1},t_{N_{\text{time}}}),z({\bf s}_{2},t_{1}),\dots,z({\bf s}_{N_{\text{space}}},t_{N_{\text{time}}})\big{)}.$
(Here we suppose $N=N_{\text{space}}N_{\text{time}}$, although any inputs ${\bf x}$ with missing temperature measurements are omitted.) The matrix ${\bf W}$ in Equation F.5 is then $N\times[(J+1)\times(1+M_{s}+M_{e}+M_{c})]$, meaning that this yields $(J+1)(1+M_{s}+M_{e}+M_{c})$ hyperparameters for the prior mean. We re-label the coefficients $\{\delta^{(\cdot)}_{jm}\}$ from Equation F.4 into an appropriately ordered vector ${\boldsymbol{\beta}}$ such that the prior mean function across all $N$ measurements can be concisely written as ${\bf W}{\boldsymbol{\beta}}$, as used in Section 3.1.
Figure F.1: Comparison of the fitted prior mean (using ordinary least squares) versus input data for four selected stations (shown in panel a.) across all time points (panel b.), where the raw data are shown in black and the fitted prior mean in red. Panel c. then shows spatial maps of the 2-, 5-, and 10-day lagged autocorrelation from the input data (top) and residuals (data minus OLS prior mean; bottom), while panel d. shows the length-scale and signal variance from semivariograms fitted separately for each for the input data (top) and residuals (data minus OLS prior mean; bottom).
After some trial and error and in order to maintain a compromise between sufficient flexibility and overfitting, we fix $J=8$, $M_{s}=200$, $M_{e}=5$, and $M_{c}=5$, resulting in $1899$ hyperparameters for the prior mean. To visualize the selected prior mean function, we show the fitted prior mean – for starters using ordinary least squares (OLS), but later these hypereparameters will be trained online with the GP – for specific stations and days versus the corresponding raw temperature measurements; see Figure F.1. Across all input locations, the OLS linear mean explains 94.1% of the variability in the daily temperature data, which is impressive since there is one hyperparameter for every $\approx 1100$ data points. For the specific stations shown in Figure F.1(b), it is clear that the seasonal cycle appropriately captures the general trend of the data, such that the spatially-varying nature of the coeffients can describe different amplitudes and even shapes in the the seasonal cycle: note that the daily maximum temperatures range from approximately 5C to 35C at the Matador station and approximately -10C to 30C at the Sibley station; furthermore, the seasonal cycle is extremely smooth at the Minden station but “bumpy” at Morehead City.
It is important to note how the variability explained by a flexible prior mean function “increases” the sparsity that can be leveraged by the kernel. In the time dimension, Figure F.1(c) shows how the 2-, 5-, and 10-day lagged autocorrelations at each gauged location differ for the input data versus the residuals (the input data minus the OLS prior mean function). Even at a lag of 10 days, the data show strong autocorrelations, in excess of 0.6. However, when removing the fitted OLS prior mean, 10-day lagged autocorrelations essentially vanish; even for 5-day lags, the autocorrelations are quite small. Spatially, Figure F.1(d) shows the length-scale and signal variance from a variogram fitted to either the input data (black) or the OLS residuals separately for each day. Here, the effect of the prior mean function is less clear: in particular, the length-scale of the residuals appear to be on a similar order of magnitude as the input data. On the other hand, the variability of the residuals is significantly reduced relative to the input data. In summary, it is clear that the prior mean function allows the data to be quite sparse in the time domain (i.e., correlations beyond ±$\pm 5$ days can be safely ignored) and perhaps less sparse in the spatial domain. These differences in the structure explained by the prior mean (temporal vs. spatial) have important implications for how we set up the kernel.
Kernel. For the full space-time data, we use a core kernel similar to the one in Equation F.1, albeit generalized to a three-dimensional input space, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})$. The anisotropy matrix again uses the eigendecomposition ${\bf\Sigma}({\bf x})={\bf\Gamma}({\bf x}){\bf\Lambda}({\bf x}){\bf\Gamma}({\bf x})^{\top}$, where
${\bf\Lambda}({\bf x})=\left[\begin{array}[]{ccc}\lambda_{1}({\bf s})&0&0\\ 0&\lambda_{2}({\bf s})&0\\ 0&0&\lambda_{3}(t)\end{array}\right],\hskip 12.91663pt{\bf\Gamma}({\bf x})=\left[\begin{array}[]{ccc}\cos\gamma({\bf s})&-\sin\gamma({\bf s})&0\\ \sin\gamma({\bf s})&\cos\gamma({\bf s})&0\\ 0&0&1\end{array}\right].$
The formulas for $\lambda_{1}({\bf s})$, $\lambda_{2}({\bf s})$, and $\gamma({\bf s})$ are as in Equation F.2, and the length-scale in time is a constant, i.e., $\lambda_{3}(t)\equiv\lambda_{3}$. Note that in this case, the term inside the Matérn correlation is
$\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}\left(\frac{{\bf\Sigma}({\bf x})+{\bf\Sigma}({\bf x}^{\prime})}{2}\right)^{-1}({\bf x}-{\bf x}^{\prime})}=\sqrt{({\bf s}-{\bf s}^{\prime})^{\top}\left(\frac{{\bf\Sigma}({\bf s})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right)^{-1}({\bf s}-{\bf s}^{\prime})+\frac{(t-t^{\prime})^{2}}{\lambda_{3}}},$
and, since
${\bf\Sigma}({\bf x})=\left[\begin{array}[]{cc}{\bf\Sigma}({\bf s})&{\bf 0}\\ {\bf 0}&\lambda_{3}\end{array}\right],$
we have
$\frac{\left|{\bf\Sigma}({\bf x})\right|^{1/4}\left|{\bf\Sigma}({\bf x}^{\prime})\right|^{1/4}}{\left|\frac{{\bf\Sigma}({\bf x}^{\prime})+{\bf\Sigma}({\bf x}^{\prime})}{2}\right|^{1/2}}=\frac{\left|{\bf\Sigma}({\bf s})\right|^{1/4}\lambda_{3}^{1/4}\left|{\bf\Sigma}({\bf s}^{\prime})\right|^{1/4}\lambda_{3}^{1/4}}{\left|\frac{{\bf\Sigma}({\bf s}^{\prime})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right|^{1/2}\lambda_{3}^{1/2}}=\frac{\left|{\bf\Sigma}({\bf s})\right|^{1/4}\left|{\bf\Sigma}({\bf s}^{\prime})\right|^{1/4}}{\left|\frac{{\bf\Sigma}({\bf s}^{\prime})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right|^{1/2}}.$
Next, the signal variance is again a log-linear function
$\log\sigma({\bf x})=\phi_{0}+\phi_{1}\text{Elev}({\bf s})+\phi_{2}\text{Dist2Coast}({\bf s}),$
which focuses on spatial heterogeneity in the signal variance. It is important to note that this specification yields a non-separable space-time kernel, i.e.,
$C_{\text{core}}({\bf x},{\bf x}^{\prime})\neq C({\bf s},{\bf s}^{\prime})\times C(t,t^{\prime}).$
For the sparse kernel, the Wendland component is again anisotropic with coordinate-specific length scales (and coordinate-specific truncation limits); following Section 4.1 we use 10 units for the spatial truncation limits and 5 days for the temporal truncation limits based on the fact that the prior mean function explains autocorrelation beyond 5 days for a large majority of the domain. We again use 100 bump functions with $n_{1}=4$ and $n_{2}=25$; however, again based on the fact that the prior mean explains so much of the temporal variability, we restrict the bump functions to live in the spatial domain only. In total, this yields $521+1+1=523$ hyperparameters for the kernel (the same 521 from the time-slice modeling in Section 4.1 plus $\lambda_{3}$ and the Wendland length-scale in the time dimension).
Appendix G Supplemental figures
Figure G.1: Sample bump functions $g({\bf x})$ from Equation 1 for different values of the shape parameter $b$ with fixed amplitude $a$ and radius $r$. As $b\rightarrow 0$, the bump function becomes uniform on $(-r,r)\subset\mathbb{R}^{1}$; as $b\rightarrow\infty$, the bump functions become delta function centered at $\mathbf{h}$. Note that the bump functions are exactly zero for distances from the centroid ${\bf h}$ greater than or equal to $r$.
Figure G.2: Posterior summaries of hyperparameters for the January 14, 2001 time slice of daily maximum temperature. Panels (a), (b), and (c) show the prior mean function, square root of the signal variance, and anisotropy length-scales, calculated using the posterior median hyperparameters and the physical covariates. Panel (d) shows implied correlations between eight reference locations and all other points in the input domain.